Ivanhoe Company issues $320,000, 20 year, 7% bonds at 101. Prepare the journal entry to record the sale of these bonds on June 1, 2022.
Answer #1
Date Accounts titles and Explanation Debit ($) Credit ($)
June 1 Cash 323,200
Bonds payable 320,000
Premium on bonds payable 3,200
(To record the sale of the bonds)

Cash - 320,000 * 101% = 323,200

Premium on bonds payable = 323,200 - 320,000 = 3,200